# MLR-Bench: Evaluating AI Agents on Open-Ended Machine Learning Research

### v2

@@ 1 Introduction
To explore RQ1, we use MLR-Bench to evaluate six state-of-the-art models—including the most recent o4-mini (o4mini2025, ), Gemini-2.5-pro-preview (gemini2025, ), and Qwen3-235B-A22B (qwenteam2025qwen3, )—as well as the advanced coding agent, Claude Code (claude2025, ). Our results show that while these models perform well in generating ideas and writing papers, they often fail to deliver innovative or scientifically reliable research, mostly due to coding agents producing invalidated experimental results.
@@ 2 MLR-Bench
In this section, we describe the overall flow of MLR-Bench. Fig. 1 presents an overview of the framework of MLR-Bench, which consists of both an end-to-end evaluation pipeline to evaluate AI agents’ ability to complete end-to-end research, and a stepwise evaluation pipeline, which separately evaluates AI agents’ abilities on (1) Idea Generation, (2) Proposal Generation, (3) Experimentation, and (4) Paper Writing.
Tasks. To measure the performance of AI agents across various machine learning topics, we collect 201 tasks from ICLR, ICML, and NeurIPS workshops over the past three years. These tasks are categorized into 9 core ML topics: LLMs and Vision Language Models (VLMs), AI for Science, ML Theory, Trustworthy AI, Computer Vision, ML Systems, Multimodality, Reinforcement Learning (RL), and other emerging topics. Fig. 2 illustrates the distribution of our tasks across these ML topics. To curate our dataset, we first review all workshops from ICLR, ICML, and NeurIPS conferences over the past three years. We then filter out duplicate workshops and select those that maintain complete information and target a general audience. Finally, we extract workshop overviews and topics to formulate our tasks.
End-to-End Evaluation. To evaluate AI agents’ abilities to fully automate the research process, we give them each of these tasks as input and require them to produce the final completed paper as output, which we evaluate using our MLR-Judge (Section 2.1). To facilitate software experimentation, agents are provided with an environment that includes file system access, a Python runtime for executing scripts, and internet connectivity.
@@
Each step of the stepwise evaluation requires a set of input data for the agents. For example, in step (1), the Idea Generation agents require research tasks as input data; we use the same 201 tasks used in end-to-end evaluation. In step (2) the Proposal Generation agents require tasks and research ideas as input. To construct this dataset, for each of the 201 tasks, we randomly sample one idea (generated by any Idea Generation agent) from step (1) for this task, resulting in 201 (task, idea) pairs that serve as the input for step (2). We apply the same procedure in steps (3) and (4), each time randomly sampling from the outputs of the previous step for a task, except in step (3) we manually select a subset of 10 (task, idea, proposal) triples based on their suitability for evaluating the experimentation agents, to keep the experiments manageable. Details about these 10 selected tasks can be found in Appendix A.
@@ 2.1 MLR-Judge
In this work, we develop an automated evaluation tool, MLR-Judge, which consists of multiple carefully designed rubrics and an LLM judge. Prior to designing these rubrics, we summarize the key review dimensions that people focus on when evaluating research outputs at each step: Consistency, Novelty, Clarity, Feasibility, Completeness, Soundness, Insightfulness, Significance and Overall. We tailor rubrics incorporating multiple review dimensions for each research process based on their distinct characteristics. Fig. 2 illustrates the review dimensions for different research processes.
Additionally, MLR-Judge incorporates an LLM judge alongside our rubrics. In this work, we select Gemini-2.5-Pro-Preview and Claude-3.7-Sonnet as our judge models, which have demonstrated superior reasoning capabilities and multimodal functionality, enabling them to review papers containing figures. These two judge models independently evaluate the quality of research outputs, and the final evaluation results are averaged from their individual assessments. At each research step, MLR-Judge evaluates the Consistency dimension by analyzing both the history information and the output of MLR-Agent. For the experimentation step, we examine the execution log of our coding agent to gain insights into the experimental process. Details of rubrics and prompts of MLR-Judge can be found in Section E.2.
@@ 2.2 MLR-Agent
MLR-Agent operates in two modes: a stepwise execution mode and an end-to-end execution mode. The stepwise execution mode enables us to evaluate each model’s performance in individual research steps, while the end-to-end execution mode allows us to assess the models’ capabilities in automating complete research projects. MLR-Agent’s automated research process follows the same research steps as illustrated in Fig. 1, namely (1) Idea Generation, (2) Proposal Generation, (3) Experimentation, and (4) Paper Writing. MLR-Agent uses LLMs for steps (1) and (2), a coding agent (Claude Code) for step (3), and multimodal LLMs for step (4). Additionally, since most frontier models currently lack web search capabilities, they face challenges in accessing recent related work. Therefore, MLR-Agent uniformly employs GPT-4o-Search-Preview (searchgpt2024, ) to conduct a literature review between (1) Idea Generation and (2) Proposal Generation, providing reference materials for the agent. Details about the prompt used for each stage can be found in Section E.1.
@@ 3 How Well Can AI Agents Conduct Open-Ended Research?
This section describes empirical studies of how well AI agents perform on open-ended research. We use MLR-Agent as our agent scaffold to conduct research and evaluate a number of frontier models, as shown in Fig. 3. In experimentation and analysis step, our coding agent executes on an Ubuntu 22.04 server with access to four NVIDIA RTX 3090 GPUs. We use MLR-Judge to evaluate the outcomes of different research processes. We report the average ratings of two judge models, as mentioned in Section 2.1, and details of each rating can be found in Appendix D.
@@
Coding Claude Code (Claude-3.7-Sonnet)
§3.3
@@ 3.1 Results of Idea Generation Evaluation
All models demonstrate strong performance in Consistency and Significance, while Feasibility and Novelty are more challenging. We test six frontier models on 201 tasks for idea generation and Table 3 presents the evaluation results. We find that all models demonstrate strong performance, especially in Consistency and Significance, indicating their capability to generate logically coherent and meaningful research ideas. However, there exists a notable performance gap in Novelty and Feasibility, suggesting that generating innovative yet implementable ideas remains a significant challenge for current LLMs.
@@ 3.2 Results of Proposal Generation Evaluation
All models excel at generating logically coherent research proposals with research significance, but novelty, soundness and feasibility are more challenging. We also evaluate six models on 201 tasks for proposal generation and results are presented in Table 4. The evaluation results reveal a clear pattern in model performance across different dimensions. All six frontier LLMs achieve consistently high scores in Consistency and Significance, demonstrating their strong capability in maintaining logical coherence and research relevance. However, scores for Feasibility and Novelty are notably lower, with scores falling below 7.5 in most cases. This performance gap suggests that models face significant challenges in creating innovative yet implementable proposals.
@@ 3.3 Results of Experimentation Evaluation
Claude Code cannot yet produce solid experimental results. We test the performance of Claude Code on 10 typical ML tasks selected from the 201 tasks. Fig. 3 shows the assessments of two LLM judge models across seven review dimensions. Although the scores of Claude-3.7-Sonnet are significantly higher than those of Gemini-2.5-Pro-Preview, both judge models reveal weaknesses in Soundness, Insightfulness, and Significance, suggesting potential limitations in the robustness of the experimental results. Additionally, neither judge model gives an overall assessment higher than 7.0, further indicating concerns about the reliability of the results.
@@ 3.4 Results of Paper Writing Evaluation
Gemini-2.5-Pro-Preview presents superior performance in paper writing. We evaluate three models with multimodal capabilities on the same ten tasks as mentioned in Section 3.3 in paper writing. The evaluation results presented in Table 5 indicate that Gemini-2.5-Pro-Preview demonstrates stronger writing capabilities, while o4-mini-high shows relatively weaker performance. From manual observation, we believe this is partly due to Gemini-2.5-Pro-Preview’s detailed formal explanations, including algorithms, formulas, and derivations, along with its strong mathematical capabilities, while o4-mini-high performs worse due to its relatively terse style affecting its clarity. However, none of the models achieve high overall scores (above 7.0), which may be influenced by the weaker experimental results from the previous step. This demonstrates that experimental success is a critical determinant of overall research quality.
@@ 3.5 Results of End-to-End Evaluation
Trade-off between performance and cost should be considered when developing research agents. On the same tasks in Section 3.4, we test three models for end-to-end automated research. The evaluation results and costs are shown in Table 6. The results demonstrate that all models achieve higher scores in Clarity and Novelty while scores of Soundness and Overall are notably lower. This indicates that current LLMs still have substantial room for improvement in end-to-end automated research. While Claude-3.7-Sonnet demonstrates superior performance across all dimensions (overall score of 4.70), it comes at a significantly higher cost (2.40 USD). Interestingly, o4-mini-high achieves competitive results (overall score of 3.95) at the lowest cost (1.15 USD), suggesting that smaller models can serve as a more economical option for research agents.
Table 6: End-to-end evaluation results and costs averaged over ten tasks.
Model Clarity Novelty Soundness Significance Overall Cost (\$)
o4-mini-high $7.65\pm 0.32$ $6.90\pm 0.30$ $3.80\pm 0.60$ $5.05\pm 0.57$ $3.95\pm 0.71$ $1.15$
Gemini-2.5-Pro-Preview $7.40\pm 0.45$ $6.75\pm 0.63$ $3.35\pm 0.88$ $4.55\pm 0.88$ $3.75\pm 0.81$ $1.24$
Claude-3.7-Sonnet $7.75\pm 0.34$ $7.10\pm 0.43$ $4.05\pm 1.11$ $5.50\pm 1.14$ $4.70\pm 1.22$ $2.40$
@@ 4 How Well Is MLR-Judge Aligned with Human Reviewers?
Agreement Score. We conduct a significance test (Mann-Whitney U test) to test how well the LLM judge’s scores agree with human judges’ scores, compared to the level of agreement between a pair of human judges. Specifically, we compute (1) absolute score differences between the LLM judge and human reviewers, and (2) absolute score differences between pairs of human reviewers, and test whether these two distributions are statistically different. As shown in Fig. 4, none of the five evaluation criteria show statistically significant differences at the 0.05 confidence level. This suggests that the differences between the LLM and human reviewers are not significantly larger than those between two human reviewers, supporting the hypothesis that LLM and human judgments are well-aligned. Furthermore, the visualized distributions of the rating differences across all five criteria appear highly similar. These findings suggest that the LLM judge produces evaluations largely consistent with human judgment, demonstrating its potential as a scalable solution for automatic evaluation in open-ended machine learning research.
@@ 6 Related Work
Recent works (radensky2024let, ; llmwritingsurvey2024, ) have explored using LLMs for various stages of scientific research, including idea generation (si2025can, ; ideabench2025, ; chainofideas2024, ), experimentation (huang2024mlagentbench, ; chan2025mlebench, ; tian2024scicode, ; miserendino2025swelancerfrontierllmsearn, ), and writing (llmwritingsurvey2024, ; xi2025omnithinkexpandingknowledgeboundaries, ; wang2024autosurvey, ). Existing benchmarks typically target narrow tasks: the MLE benchmark (chan2025mlebench, ) focuses on engineering, the MLAgentBench huang2024mlagentbench on experimentation, and the PaperBench starace2025paperbenchevaluatingaisability on paper reproduction, each revealing gaps between AI agents and human researchers (chan2025mlebench, ; huang2024mlagentbench, ; starace2025paperbenchevaluatingaisability, ). RE-Bench rebench2025 pushes toward generalization on unseen tasks (rebench2025, ). Our proposed MLR-Bench complements these efforts by evaluating 201 open-ended research tasks from major Machine Learning conferences, covering the entire research pipeline. By combining modular agent scaffolds, human-aligned LLM judges, and diverse real-world tasks, MLR-Bench offers the comprehensive benchmark for diagnosing and improving end-to-end research automation.

### v3

@@ 1 Introduction
To explore RQ1, we use MLR-Bench to evaluate six state-of-the-art models—including the most recent o4-mini (o4mini2025, ), Gemini-2.5-pro-preview (gemini2025, ), and Qwen3-235B-A22B (qwenteam2025qwen3, )—as well as the advanced coding agents, Claude Code (claude2025, ) and Codex (codex2025, ). Our results show that while these models perform well in generating ideas and writing papers, they often fail to deliver innovative or scientifically reliable research, mostly due to coding agents producing invalidated experimental results.
@@ 2 MLR-Bench
In this section, we describe the overall flow of MLR-Bench. Fig.˜1 presents an overview of the framework of MLR-Bench, which consists of both an end-to-end evaluation pipeline to evaluate AI agents’ ability to complete end-to-end research, and a stepwise evaluation pipeline, which separately evaluates AI agents’ abilities on (1) Idea Generation, (2) Proposal Generation, (3) Experimentation, and (4) Paper Writing.
Tasks. To measure the performance of AI agents across various machine learning topics, we collect 201 tasks from ICLR, ICML, and NeurIPS workshops over the past three years. These tasks are categorized into 9 core ML topics: LLMs and Vision Language Models (VLMs), AI for Science, ML Theory, Trustworthy AI, Computer Vision, ML Systems, Multimodality, Reinforcement Learning (RL), and other emerging topics. Fig.˜2 illustrates the distribution of our tasks across these ML topics. To curate our dataset, we first review all workshops from ICLR, ICML, and NeurIPS conferences over the past three years. We then filter out duplicate workshops and select those that maintain complete information and target a general audience. Finally, we extract workshop overviews and topics to formulate our tasks.
End-to-End Evaluation. To evaluate AI agents’ abilities to fully automate the research process, we give them each of these tasks as input and require them to produce the final completed paper as output, which we evaluate using our MLR-Judge (Section˜2.1). To facilitate software experimentation, agents are provided with an environment that includes file system access, a Python runtime for executing scripts, and internet connectivity.
@@
Each step of the stepwise evaluation requires a set of input data for the agents. For example, in step (1), the Idea Generation agents require research tasks as input data; we use the same 201 tasks used in end-to-end evaluation. In step (2) the Proposal Generation agents require tasks and research ideas as input. To construct this dataset, for each of the 201 tasks, we randomly sample one idea (generated by any Idea Generation agent) from step (1) for this task, resulting in 201 (task, idea) pairs that serve as the input for step (2). We apply the same procedure in steps (3) and (4), each time randomly sampling from the outputs of the previous step for a task, except in step (3) we manually select a subset of 10 (task, idea, proposal) triples based on their suitability for evaluating the experimentation agents, to keep the experiments manageable. Details about these 10 selected tasks can be found in Appendix˜A.
@@ 2.1 MLR-Judge
In this work, we develop an automated evaluation tool, MLR-Judge, which consists of multiple carefully designed rubrics and an LLM judge. Prior to designing these rubrics, we summarize the key review dimensions that people focus on when evaluating research outputs at each step: Consistency, Novelty, Clarity, Feasibility, Completeness, Soundness, Insightfulness, Significance and Overall. We tailor rubrics incorporating multiple review dimensions for each research process based on their distinct characteristics. Fig.˜2 illustrates the review dimensions for different research processes.
Additionally, MLR-Judge incorporates an LLM judge alongside our rubrics. In this work, we select Gemini-2.5-Pro-Preview and Claude-3.7-Sonnet as our judge models, which have demonstrated superior reasoning capabilities and multimodal functionality, enabling them to review papers containing figures. These two judge models independently evaluate the quality of research outputs, and the final evaluation results are averaged from their individual assessments. At each research step, MLR-Judge evaluates the Consistency dimension by analyzing both the history information and the output of MLR-Agent. For the experimentation step, we examine the execution log of our coding agent to gain insights into the experimental process. Details of rubrics and prompts of MLR-Judge can be found in LABEL:app-subsec:judge-prompts.
@@ 2.2 MLR-Agent
MLR-Agent operates in two modes: a stepwise execution mode and an end-to-end execution mode. The stepwise execution mode enables us to evaluate each model’s performance in individual research steps, while the end-to-end execution mode allows us to assess the models’ capabilities in automating complete research projects. MLR-Agent’s automated research process follows the same research steps as illustrated in Fig.˜1, namely (1) Idea Generation, (2) Proposal Generation, (3) Experimentation, and (4) Paper Writing. MLR-Agent uses LLMs for steps (1) and (2), a coding agent (e.g., Claude Code) for step (3), and multimodal LLMs for step (4). Additionally, since most frontier models currently lack web search capabilities, they face challenges in accessing recent related work. Therefore, MLR-Agent uniformly employs GPT-4o-Search-Preview (searchgpt2024, ) to conduct a literature review between (1) Idea Generation and (2) Proposal Generation, providing reference materials for the agent. Details about the prompt used for each stage can be found in Section˜D.1.
@@ 3 How Well Can AI Agents Conduct Open-Ended Research?
This section describes empirical studies of how well AI agents perform on open-ended research. We use MLR-Agent as our agent scaffold to conduct research and evaluate a number of frontier models, as shown in Fig.˜3. In experimentation and analysis step, our coding agents execute on an Ubuntu 22.04 server with access to four NVIDIA RTX 3090 GPUs. We use MLR-Judge to evaluate the outcomes of different research processes. We report the average ratings of two judge models, as mentioned in Section˜2.1, and details of each rating can be found in Appendix˜C.
Table 2: Evaluated models (o4mini2025, ; claude2025, ; deepseekai2025deepseekr1incentivizingreasoningcapability, ; ministral2024, ; qwenteam2025qwen3, ; gemini2025, ; codex2025, ; geminicli2025, ) in different research stages.
@@
Coding Claude Code (Claude-3.7-Sonnet),
§3.3 Codex (o4-mini)
@@ 3.1 Results of Idea Generation Evaluation
All models demonstrate strong performance in Consistency and Significance, while Feasibility and Novelty are more challenging. We test six frontier models on 201 tasks for idea generation and Table˜3 presents the evaluation results. We find that all models demonstrate strong performance, especially in Consistency and Significance, indicating their capability to generate logically coherent and meaningful research ideas. However, there exists a notable performance gap in Novelty and Feasibility, suggesting that generating innovative yet implementable ideas remains a significant challenge for current LLMs.
@@ 3.2 Results of Proposal Generation Evaluation
All models excel at generating logically coherent research proposals with research significance, but novelty, soundness and feasibility are more challenging. We also evaluate six models on 201 tasks for proposal generation and results are presented in Table˜4. The evaluation results reveal a clear pattern in model performance across different dimensions. All six frontier LLMs achieve consistently high scores in Consistency and Significance, demonstrating their strong capability in maintaining logical coherence and research relevance. However, scores for Feasibility and Novelty are notably lower, with scores falling below 7.5 in most cases. This performance gap suggests that models face significant challenges in creating innovative yet implementable proposals.
@@ 3.3 Results of Experimentation Evaluation
Popular coding agents cannot yet produce solid experimental results. We evaluate the experimental execution by comparing Claude Code and Codex on 10 representative machine learning tasks selected from the 201 tasks. Our LLM judge assesses both the execution records and the final results. As presented in Table˜5, the overall ratings for both agents fall below the 6.0 acceptance threshold, indicating that current coding agents still struggle to generate robust experimental outcomes. Specifically, while Claude Code achieves respectable scores for Consistency, Completeness, and Novelty, it falters on Soundness, Insightfulness, and Significance. This weakness is detailed in Table˜5 and further highlighted in Fig.˜3, where the Gemini judge’s scores for its Soundness and Significance are below 3.0. This suggests that although Claude Code can devise and execute complete and novel experiments, the results often lack scientific robustness. In contrast, Codex performs well on Soundness but underperforms on Novelty, Insightfulness, and Significance. This suggests that while Codex is reliable in executing experiments, it lacks the capability to design novel ones.
Table 5: Evaluation results of Claude Code and Codex averaged on ten tasks in experiment execution. The scores are the average of two judge models.
Coding Agent Consistency Completeness Novelty Soundness Insightfulness Significance Overall
Claude Code $6.75\pm 1.00$ $6.00\pm 0.68$ $5.65\pm 0.82$ $4.75\pm 1.02$ $4.50\pm 0.97$ $4.70\pm 0.95$ $4.95\pm 0.82$
Codex $6.30\pm 1.46$ $5.05\pm 0.87$ $3.80\pm 1.19$ $6.15\pm 0.87$ $4.45\pm 1.28$ $3.40\pm 1.05$ $4.95\pm 1.02$
@@ 3.4 Results of Paper Writing Evaluation
Gemini-2.5-Pro-Preview presents superior performance in paper writing. We evaluate three models with multimodal capabilities on the same ten tasks as mentioned in Section˜3.3 in paper writing. The evaluation results presented in Table˜6 indicate that Gemini-2.5-Pro-Preview demonstrates stronger writing capabilities, while o4-mini-high shows relatively weaker performance. From manual observation, we believe this is partly due to Gemini-2.5-Pro-Preview’s detailed formal explanations, including algorithms, formulas, and derivations, along with its strong mathematical capabilities, while o4-mini-high performs worse due to its relatively terse style affecting its clarity. However, none of the models achieve high overall scores (above 7.0), which may be influenced by the weaker experimental results from the previous step. This demonstrates that experimental success is a critical determinant of overall research quality.
@@ 3.5 Results of End-to-End Evaluation
Current research agents struggle to generate robust scientific results.
We conduct an end-to-end evaluation to compare three reasoning models within our MLR-Agent scaffold and benchmark them against AI Scientist V2 (Yamada2025ai, ), with results summarized in Table˜7. Our findings reveal two key points. Firstly, when using the MLR-Agent scaffold, the overall ratings for all three models fall below the 6.0 acceptance threshold. Notably, all models receive their lowest scores on Soundness, which underscores their shared difficulty in producing scientifically sound results. Secondly, in a direct model comparison, o4-mini is significantly outperformed by Gemini and Claude, whose performances are comparable. This trend is inversely correlated with cost. When using Claude Code as the coding agent, o4-mini is the most affordable, followed by Gemini. Consequently, Gemini emerges as the most cost-effective model for end-to-end autonomous research, considering its strong performance and moderate cost.
Table 7: End-to-end evaluation results over the ten tasks. AI Scientist V2 is powered by o4-mini-high. We use o4-mini-medium in Codex and use Gemini-2.5-Pro-Preview-05-06 in this experiment. When using Claude Code as the coding agent, the costs of o4-mini-high, Gemini-2.5-Pro-Preview, and Claude-3.7-Sonnet are \$1.15, \$1.24, and \$2.40, respectively.
Model Clarity Novelty Soundness Significance Overall
AI Scientist V2
o4-mini-high $6.55\pm 0.94$ $6.70\pm 0.48$ $3.70\pm 1.29$ $4.85\pm 1.08$ $4.25\pm 1.25$
MLR-Agent
o4-mini-high + Codex $6.45\pm 0.90$ $5.65\pm 0.60$ $2.90\pm 0.57$ $3.80\pm 0.65$ $3.10\pm 0.60$
Gemini-2.5-Pro-Preview + Gemini CLI $8.30\pm 0.37$ $6.85\pm 0.34$ $4.15\pm 1.06$ $5.30\pm 0.91$ $4.60\pm 1.00$
Claude-3.7-Sonnet + Claude Code $7.75\pm 0.34$ $7.10\pm 0.43$ $4.05\pm 1.11$ $5.50\pm 1.14$ $4.70\pm 1.22$
AI Scientist V2 outperforms MLR-Agent, yet both share fundamental limitations.
We further compare MLR-Agent with AI Scientist V2 in Table˜7. Our comparison between MLR-Agent and AI Scientist V2, both using o4-mini as the backbone model, reveals that AI Scientist V2 consistently outperforms MLR-Agent across all review dimensions. However, neither agent meets the 6.0 acceptance threshold. A closer analysis of their detailed scores highlights three key findings:
  • Both agents excel at generating novel research outcome clearly. Both agents demonstrate strong creativity, reflected in high scores for Novelty. They can propose original research questions crucial for scientific discovery and articulate these ideas clearly, as shown by their strong performance in the Clarity dimension.
  • Agents struggle with technical soundness. A primary weakness is their inability to propose technically sound methods. This is evident from their markedly low scores in Soundness, which consistently fall below the acceptance threshold. This suggests that while the agents can formulate creative ideas, their technical execution is frequently flawed.
  • The agents’ research demonstrates limited potential impact. This limitation is reflected in their consistently low scores for Significance, suggesting their work often fails to address important research questions. This perceived lack of contribution, potentially compounded by reproducibility challenges, severely restricts the research’s influence and uptake by the scientific community.
@@ 4 How Well Is MLR-Judge Aligned with Human Reviewers?
Agreement Score. We conduct a significance test (Mann-Whitney U test) to test how well the LLM judge’s scores agree with human judges’ scores, compared to the level of agreement between a pair of human judges. Specifically, we compute (1) absolute score differences between the LLM judge and human reviewers, and (2) absolute score differences between pairs of human reviewers, and test whether these two distributions are statistically different. As shown in Fig.˜4, none of the five evaluation criteria show statistically significant differences at the 0.05 confidence level. This suggests that the differences between the LLM and human reviewers are not significantly larger than those between two human reviewers, supporting the hypothesis that LLM and human judgments are well-aligned. Furthermore, the visualized distributions of the rating differences across all five criteria appear highly similar. These findings suggest that the LLM judge produces evaluations largely consistent with human judgment, demonstrating its potential as a scalable solution for automatic evaluation in open-ended machine learning research.
@@ 6 Related Work
Recent works (radensky2024let, ; llmwritingsurvey2024, ; zhao2025sciarena, ) have explored using LLMs for various stages of scientific research, including idea generation (si2025can, ; ideabench2025, ; chainofideas2024, ), experimentation (huang2024mlagentbench, ; chan2025mlebench, ; tian2024scicode, ; miserendino2025swelancerfrontierllmsearn, ), and writing (llmwritingsurvey2024, ; xi2025omnithinkexpandingknowledgeboundaries, ; wang2024autosurvey, ). Existing benchmarks typically target narrow tasks: the MLE benchmark (chan2025mlebench, ) focuses on engineering, the MLAgentBench huang2024mlagentbench on experimentation, and the PaperBench starace2025paperbenchevaluatingaisability on paper reproduction, each revealing gaps between AI agents and human researchers (chan2025mlebench, ; huang2024mlagentbench, ; starace2025paperbenchevaluatingaisability, ). RE-Bench rebench2025 pushes toward generalization on unseen tasks (rebench2025, ). Our proposed MLR-Bench complements these efforts by evaluating 201 open-ended research tasks from major Machine Learning conferences, covering the entire research pipeline. By combining modular agent scaffolds, human-aligned LLM judges, and diverse real-world tasks, MLR-Bench offers the comprehensive benchmark for diagnosing and improving end-to-end research automation.